OpenStudy (dan815):

r=f(t) r''=K/r^2 solve for f(t)

OpenStudy (zzr0ck3r):

f(t) = r

OpenStudy (zzr0ck3r):

:)

OpenStudy (dan815):

haha =[

OpenStudy (loser66):

hahahah... 100 for @zzr0ck3r

OpenStudy (dan815):

how about i solve it for small r increments, it wud be approximately linear ode there

OpenStudy (anonymous):

What if you multiply by the first derivative: \[\begin{align*} \color{red}{r'}r''&=\frac{K\color{red}{r'}}{r^2} \end{align*}\] To get this equation, we would have differentiated the following with respect to \(t\): \[\begin{align*} \frac{1}{2}(r')^2&=C-\frac{K}{r} \end{align*}\] I suppose there's nothing stopping us from taking the square root, then we have \[\begin{align*} r'&=\pm\sqrt{2C-\frac{2K}{r}} \end{align*}\] I'm curious to see if there's a way to find \(r\) using a similar approach for reducing \(r''\) to \(r'\)...

OpenStudy (dan815):

hey thats very nice at least we can show that r' is converging now

OpenStudy (dan815):

the problem came out of a particle charge that is being repelled by a stationary charge